solve the following using the standard identity (a+b)(a-b)=a2-b2
104x96
 

Grade:8

2 Answers

Arun
6 years ago
Dear student
 
We can write it as
104 X 96 = ( 100 + 4) X (100 -4)
This is the same form of (a+b) (a - b)
Which equals a² - b²
Hence
104 X 96 = (100 +4) X (100 - 4) = 100² - 4² = 10000 -16
 = 9984
 
Regards
Arun (askIITians forum expert)
Sayantan Garai
5 years ago
104×96
=(100+4)(100-4)........(i)
Given eq.is (a+b)(a-b)....(ii)
Comparing,we get
a=100 and b=4
Hence, 96×104=(100+4)(100-4)=100²-4²=10000-16=9984.